A private helicopter flies 250 km west from city A to city B, then covers 100 km north from city B to city C. How far is it from city A to city C? What direction must the helicopter head to return straitly to city A from city C?


1
Expert's answer
2021-11-07T19:26:29-0500

The distance between A and C and the direction to get back to A from C are


"D_{AC}=\sqrt{250^2+100^2}=269\text{ km},\\\space\\
\theta=\arctan\frac{100}{250}=21.8°\text{ South of East}."